OpenStudy (anonymous):

Suppose y partly varies directly as x^(2) and partly varies directly as x. If x=1, then y=14; if x=6, then y=144. (a) Express y in terms of x. => y=2x^(2)+12x (b) Find the values of x when y=32. => x=2/-8 (c) Find the minimum value of y. i am doubted in c...

OpenStudy (amistre64):

pertly varies? what does that mean ....

OpenStudy (amistre64):

if a and b are correct ... then the minimum point of a parabola is its vertex

OpenStudy (anonymous):

i forgot how to calculate the vertex

OpenStudy (amistre64):

y = 2x^2 + 12x y = 2(x^2 + 6x) y = 2(x^2 + 6x +9 -9) y = 2((x+3)^2 -9) y = 2(x+3)^2 -18

OpenStudy (amistre64):

you can either calculate it, or restructure the equation into its vertex form

OpenStudy (anonymous):

ohh thanks

OpenStudy (amistre64):

the last term, what would normally be a y intercept ... in a quadratic equation is the y value of the vertex

OpenStudy (amistre64):

so the minimum is ... ??

OpenStudy (anonymous):

-18

OpenStudy (amistre64):

-18 is correct

OpenStudy (anonymous):

thank you very much:)